OpenStudy (anonymous):

Suppose the cost of manufacturing 10 units of a product is R40 and the cost of 20 units is R70. If the cost c is linearly related to output Q (units produced). Find the cost of producing 35 items. [1] R115,00 [2] R121,67 [3] R113,33 [4] R65,00 [5] None of the above.

OpenStudy (anonymous):

Form an equation. Gradient = (70 - 40) / (20 - 10) = 30 / 10 = 3 Using (10 , 40): c = 40 + 3(Q - 10) = 3Q + 10 Q = 35 units c = 3(35) + 10 = R115

OpenStudy (anonymous):

Hello Turitw Thank you so much for this answer. I'm sorry, my internet connection dropped last night. I really do appreciate it. It helped a lot